OpenStudy (anonymous):

How do I simplify (x^2 - 6x + 8/2x^2 - 7x +6) - (x^2 - x - 12/ 2x^2 - 5x - 12)?

OpenStudy (anonymous):

x^2 - 6x + 8 x^2 - x - 12 ----------- - ----------- 2x^2 - 7x +6 2x^2- 5x - 12 = (x-4)(x-2) (x-4)(x+3) --------- - ---------- (2x-3)(x-2) (2x+3)(x-4)

OpenStudy (anonymous):

go ahead u r right !!!!!

OpenStudy (anonymous):

= (x-4)(2x+3) (x+3)(2x-3) ---------- - ------------ (2x-3)(2x+3) (2x+3)(2x-3) = (x-4)(2x+3) - (x+3)(2x-3) ----------------------- (2x-3)(2x+3) = 2x^2 + 5x - 12 - 2x^2 - 3x +9 ---------------------------- (2x-3)(2x+3) = -8x - 3 ------- (2x-3)(2x+3)

OpenStudy (anonymous):

is the answer -8 - 3 the right answer? :( --------- (2x-3)(2x+3)

OpenStudy (anonymous):

no recheck it

OpenStudy (anonymous):

what's wrong?

OpenStudy (anonymous):

\[\frac{2x^2 -5x -12 - 2x^2 - 3x + 9}{4x^2 - 9} = \frac{-(2x+3)}{(2x-3)(2x+3)} = \frac{-1}{2x-3}\]

OpenStudy (anonymous):

-5x was wrong

OpenStudy (anonymous):

I've asked two people here and both of you gave me two different answers D:

OpenStudy (anonymous):

hahahaha which one is correct??????????

OpenStudy (anonymous):

no reply milliex tell me is it correct or not ?

OpenStudy (anonymous):

I don't know D:

OpenStudy (anonymous):

because one had -7/(2x-3)(2x+3), you have -1/2x-3 and i have -8x-3/(2x-3)(2x+3)

OpenStudy (anonymous):

upto my knowledge i m right